# Med-R2: Crafting Trustworthy LLM Physicians through Retrieval and Reasoning of Evidence-Based Medicine

### v1

@@ 4.2 Baselines
We compare Med-R2 with the following baselines: (1) The simplest baseline is Direct Response, where the model answer medical questions directly without the aid of external knowledge bases or dataset fine-tuning. (2) Vanilla RAG Lewis et al. (2020) utilizes raw queries for evidence searching, and the retrieved documents are then directly integrated into the generation process without any further manipulation. (3) Fine-Tuning leverages medical datasets to further train the model under supervised conditions. Here we have employed two distinct strategies: within-dataset fine-tuning, where the datasets for training and evaluation are derived from different parts of the same data corpus, and cross-dataset fine-tuning, where the model is fine-tuned on one medical dataset (e.g., MedMCQA) and then evaluated on different datasets (e.g., PubMedQA). (4) LLM-AMT Wang et al. (2024) is a dedicated process tailored for biomedical question answering, which includes typical modules such as query augmenter, hybrid retriever, knowledge refiner, etc. Details are discussed in Section E.2.
@@ 4.3 Main Results
We conduct evaluations to validate the efficiency of Med-R2 across different open-source models. We summarize the observations below.
Med-R2 is effective across different models. Table 2 shows that the incorporation of external knowledge bases significantly enhances the model’s ability to address medical queries, where even the most basic vanilla RAG method depicts an average enhancement of 17.92% over direct responses. Furthermore, Med-R2 provides an added layer of the enhancement by adhering to the EBM process, which outperforms all baselines across various benchmarks, achieving an average improvement of 35.46% over the direct response strategy. It is Notably, for lightweight models such as Qwen-2.5-7B and LLaMA3.1-8B, Med-R2 demonstrates increases of 79.72% and 77.80% respectively. We surmise that this is due to the fact that while lightweight models inherently lack comprehensive domain-specific medical knowledge, they possess the capability to efficiently read and identify information from external medical documents. Consequently, effective augmentation from external knowledge bases substantially bolsters the models’ capacity to tackle medical domain questions. We also ablate components of Med-R2 in Table 8 to further identify the contribution of each module.
Med-R2 shows superiority compared to fine-tuning. From Table 2, we find that Med-R2 stands out as the only approach among those leveraging external knowledge bases that surpasses the average performance of fine-tuning methods. Specifically, Med-R2 exhibits nearly equivalent performance to fine-tuning strategies in within-dataset training, yet it significantly outperforms in cross-dataset training, achieving an enhancement of 7.02% and an overall capability improvement of 3.59%. One contributing factor is that during within-dataset fine-tuning, the training and testing datasets are of the same origin, thus a model trained on homogeneous data should logically achieve substantial performance gains on the test set. Conversely, in cross-dataset fine-tuning, the heterogeneity between the training and testing datasets more rigorously assesses the model’s ability to generalize within the same domain. Under these circumstances, the utilization of a comprehensive external medical knowledge base and the effective retrieval and extraction of pertinent information becomes particularly crucial. Additionally, considering that fine-tuning necessitates additional training time and computational resources, Med-R2 emerges as a more efficient approach for enhancing the model’s performance on medical domain-related issues.
@@ 5 Ablations and Analysis Across Scales
Integrating CoT at the onset may bring adverse effects. We ablate the components of query reformulator $q_{EBM}\left[+q_{CoT}\right]$ through integrating the chain-of-thought (CoT) sequences generated by models into the initial evidence retrieval phase, which is denoted as Med-R2-CoT. Figure 4 presents a grouped bar chart illustrating the medical task performance of Med-R2-CoT and all baselines included in our main experiments. It is observed that Med-R2-CoT exhibits a decline in performance compared to Med-R2, with the disparity increasing as the model parameter scale decreases. Notably, at the 7-8 billion parameter level (e.g., Qwen2.5-7B and LLaMA3.1-8B), the performance of Med-R2-CoT is even inferior to that of the vanilla RAG strategy. We hypothesize that one contributing factor is that the lightweight models’ less solid grasp of medical knowledge. Consequently, without the aid of an external medical knowledge base, these models are unable to think in the direction of the original query, and thus, the generated CoT sequences may even negatively impact the retrieval effectiveness. However, this phenomenon is somewhat mitigated as the model parameter scale increases.
@@
6 Conclusion
Despite the promising applications of Large Language Models (LLMs) in the medical field, the highly professional nature of medical knowledge poses several challenges, including inefficiency in knowledge acquisition, low precision in medical evidence retrieval, and low effectiveness in key information extraction. In this study, we follow the Evidence-Based Medicine (EBM) process to design a novel LLM physician framework, which effectively leverages the retrieval, filtering, and reasoning processes inherent to EBM, thereby significantly enhancing the model’s performance in medical scenarios. Comprehensive experiments demonstrate that our Med-R2 holds superior advantages over existing strategies, particularly in achieving performance nearly on par with models trained on corresponding datasets, while also reducing the substantial computational costs associated with model training. Furthermore, our analysis of model scale and context window size also highlights the scaling capabilities of Med-R2.
7 Limitations
There are some limitations in our work. Firstly, due to the professional nature and rapid evolution of medical knowledge, the medical knowledge corpus we construct may not ensure complete coverage of medical information, potentially limiting the upper bound of retrieval precision. Additionally, the classification framework of our query classifier may not be comprehensive in scope, and since the classifier relies on advanced language models, it cannot guarantee the absolute accuracy in classification. Concurrently, during the fine-grained reranking phase, to balance the computational cost and effectiveness, we employ a lightweight proxy model to calculate the usefulness score of the current evidence document, yet it does not fully represent the performance tendencies of the target model during actual evaluating.
@@ Appendix A Evidence-Based Medicine (EBM)
Evidence-Based Medicine (EBM) is defined as the conscientious, explicit, and judicious application of the best current evidence in making decisions regarding the care of individual patients. This practice entails the integration of individual clinical expertise with the most reliable external clinical evidence derived from systematic research Sackett et al. (1996). The EBM process typically includes five stages, question formulating, evidence searching, evidence appraising, evidence applying and effect assessing, which aims to make the best possible health care decision through iterative improvements.
To utilize "best evidence”, researchers assess the quality of trials by determining the grading system based on the likelihood that the methods used and the results obtained are less prone to bias and more reliable. The hierarchy of evidence guides the clinical decision-making, since not all evidence is created equal, as described in Figure 6. The evidence hierarchy establishes the priority of references, particularly when conflicting facts are present within the retrieved evidence.
@@
  • Cohort Studies fall into the third-highest category of evidence. Both retrospective and prospective cohort studies are prone to various biases. Prospective cohort studies are considered more reliable, less susceptible to information biases (selection, misclassification, recall), and can establish temporal associations (outcomes following exposure). However, both types of cohort studies may suffer from confounding biases, which is a major concern that can undermine the validity of their findings.
@@
  • Expert Opinion is considered the lowest level of evidence due to its high susceptibility to bias. Compared to other levels, experts are more likely to selectively choose evidence that confirms their preconceived hypotheses or beliefs, potentially leading to conflicts of interest and a focus on a specific domain while overlooking broader contexts, thereby introducing bias into their perspectives.
Appendix B Question Formulation Details
B.1 Details of Query Classification
@@
Evidence-Based Medicine (EBM) Categories  Non-professional queries may fail to clearly articulate the current medical symptoms, thereby hindering the effective retrieval of the necessary medical evidence. Here we classify the medical queries into several types according to the categories of Evidence-Based Medicine (EBM) questions, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc., and then conduct and professional medical query reformulations based on their EBM categorization to emphasize specialized retrieval. Instructions for targeted augmentation of clinical queries are outlined in Table 4, where we employ Qwen2.5-72B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as our question reformulator.
@@
B.2 Details of Query Reformulator
@@
Following the instructions outlined in Table 4, we have employed Qwen2.5-72B-Instruct to professionally reformulate the queries within the medical datasets to enhance the subsequent retrieval of relevant evidence pertaining to the current query from the knowledge corpus.
Appendix C Evidence Reranking Details
@@
We detail the scoring implementation for the fine-grained reranking phase of retrieved documents, which is based on the categorization and mapping to general document categories, as depicted in Algorithm 1.
@@
Appendix D Evidence Assessment Details
Here, we provide the details of the iterative multi-document retrieval process, as illustrated in Algorithm 2.
@@ Appendix D Evidence Assessment Details
Appendix E Experiments Details
E.1 Details of Medical Datasets
@@
  • MedQA-USMLE and MedQA-MCMLE Jin et al. (2020): The MedQA dataset is derived from the professional medical board exams, covering languages of English, simplified Chinese, and traditional Chinese. In this study, we employ the subsets collected from US and Mainland China. The MedQA-USMLE dataset includes text materials from 18 English medical textbooks, while the MedQA-MCMLE dataset is constructed from 33 simplified Chinese medical textbooks.
  • MedMCQA Pal et al. (2022): The MedMCQA dataset is a large-scale multi-choice question answering data corpus that covers 2,400 healthcare topics and 21 medical subjects. Topics of MedMCQA span from medicine (endocrinology, infection, haematology, respiratory, etc.), surgery (general surgery, endocrinology, breast, and vascular surgery, etc.) to radiology & biochemistry, which are sourced from both real-world scenarios and simulated examinations.
  • PubMedQA Jin et al. (2019): The PubMedQA dataset is a novel biomedical question answering dataset originated from PubMed abstracts. To ensure equity, we exclude the origin-given contexts during the retrieval phase. Here we choose the 1,000 expert-annotated instances for downstream task evaluation. PubMedQA is included only in our test set due to the limited number of instances.
  • MMLU-Med Hendrycks et al. (2021): MMLU-Med is a set of six medical tasks (anatomy, clinical knowledge, professional medicine, human genetics, college medicine, college biology) derived from the MMLU dataset. MMLU-Med is also included only in our test set due to its limited quanity.
@@
MedQA-USMLE Jin et al. (2020) 10178 2535 10 12723
MedQA-MCMLE Jin et al. (2020) 27400 6840 11 34251
MedMCQA Pal et al. (2022) 182822 4170 13 187005
PubMedQA Jin et al. (2019) - 990 10 1000
MMLU-Med Hendrycks et al. (2021) - 1080 9 1089
E.2 Details of Baselines
  • Direct Response: In this instance, we employ the base model to directly respond to medical queries without the aid of additional training on any dataset or augmentation from external knowledge bases.
  • Vanilla RAG: It represents the most traditional and fundamental strategy for utilizing external knowledge bases to assist models in answering questions. In this study, we employ the same medical knowledge base as Med-R2 , but directly utilizing the raw text retrieved by FAISS and combining it with the original query to test the model’s performance on medical tasks.

### v4

@@ 4.2 Baselines
We compare Med-R2 with the following baselines: (1) Frontier Models contain GPT-4o [17], Claude3.5-Sonnet [3] and DeepSeek-V3 [30]. (2) Open-Sourced Medical Models include PMC-LLaMA-7B, PMC-LLaMA-13B [47], MEDITRON-7B and MEDITRON-70B [7]. (3) The simplest baseline is Direct Response, where the model answer medical questions directly without the aid of external knowledge bases or dataset fine-tuning. (4) Vanilla RAG [28] utilizes raw queries for evidence searching, and the retrieved documents are then directly integrated into the generation process without any further manipulation. (5) Fine-Tuning leverages medical datasets to further train the model under supervised conditions. Here we have employed two distinct strategies: within-dataset fine-tuning, where the datasets for training and evaluation are derived from different parts of the same data corpus, and cross-dataset fine-tuning, where the model is fine-tuned on one medical dataset (e.g., MedMCQA) and then evaluated on different datasets (e.g., PubMedQA). (6) LLM-AMT [46] is a dedicated process tailored for biomedical question answering, which includes typical modules such as query augmenter, hybrid retriever, knowledge refiner, etc. Details are discussed in Section C.2.
@@ 4.3 Main Results
We conduct evaluations to validate the efficiency of Med-R2 across different open-sourced models. We summarize the observations below.
Med-R2 is effective across different models. Table 2 shows that the incorporation of external knowledge bases significantly enhances the model’s ability to address medical queries, where even the most basic vanilla RAG method depicts an average enhancement of 17.92% over direct responses. Furthermore, Med-R2 provides an added layer of the enhancement by adhering to the EBM process, which outperforms all baselines across various benchmarks, achieving an average improvement of 35.10% over the direct response strategy. Notably, for lightweight models such as Qwen-2.5-7B and LLaMA3.1-8B, Med-R2 demonstrates increases of 79.16% and 77.06% respectively. We surmise that this is due to the fact that while lightweight models inherently lack comprehensive domain-specific medical knowledge, they possess the capability to efficiently read and identify information from external medical documents. Consequently, effective augmentation from external knowledge bases substantially bolsters the models’ capacity to tackle medical-domain questions. Compared with medical-specific models of the same parameter scale, models equipped with our Med-R2 demonstrate superior performance in medical tasks. Moreover, LLaMA3.1-70B + Med-R2 has the potential to surpass frontier models under medical scenarios, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80%. In order to evaluate Med-R2 in real clinical settings, we have performed experiments on several datasets related to real-world clinical scenarios as well, detailed in Section C.4. Additionally, we have also ablate the components of Med-R2  in Table 10 to further identify the contribution of each module.
Med-R2 shows superiority compared to fine-tuning. From Table 2, we find that Med-R2 stands out as the only approach among those leveraging external knowledge bases that surpasses the average performance of fine-tuning methods. Specifically, Med-R2 exhibits nearly equivalent performance to fine-tuning strategies in within-dataset training, yet it significantly outperforms in cross-dataset training, achieving an enhancement of 8.59% and an overall capability improvement of 3.32%. One contributing factor is that during within-dataset fine-tuning, the training and testing datasets are of the same origin, thus a model trained on homogeneous data should logically achieve substantial performance gains on the test set. Conversely, in cross-dataset fine-tuning, the heterogeneity between the training and testing datasets more rigorously assesses the model’s ability to generalize within the same domain. Under these circumstances, the utilization of a comprehensive external medical knowledge base and the effective retrieval and extraction of pertinent information becomes particularly crucial. This phenomenon is even more pronounced in Table 8 when evaluating models in real clinical scenarios where the complexity and variability of patient cases require a higher degree of information retrieval and evidence synthesis. Additionally, considering that fine-tuning necessitates additional training time and computational resources, Med-R2 emerges as a more efficient approach for enhancing the model’s performance on medical domain-related issues.
@@ 5 Ablations and Analysis Across Scales
Integrating CoT at the onset may bring adverse effects. We ablate the components of query reformulator $q_{EBM}\left[+q_{CoT}\right]$ through integrating the chain-of-thought (CoT) sequences generated by models into the initial evidence retrieval phase, which is denoted as Med-R2-CoT. Figure 4 presents a grouped bar chart illustrating the medical task performance of Med-R2-CoT and all baselines included in our main experiments. It is observed that Med-R2-CoT exhibits a decline in performance compared to Med-R2, with the disparity increasing as the model parameter scale decreases. Notably, at the 7-8 billion parameter level (e.g., Qwen2.5-7B and LLaMA3.1-8B), the performance of Med-R2-CoT is even inferior to that of the vanilla RAG strategy. We hypothesize that one contributing factor is that the lightweight models’ less solid grasp of medical knowledge. As a result, in the absence of supplementary external medical knowledge bases, models of these scales may struggle to align the direction of their thinking with the original query, and thus, the generated CoT sequences may even negatively impact the retrieval effectiveness. However, this phenomenon is somewhat mitigated as the model parameter scale increases.
@@
6 Discussion
In this study, we follow the Evidence-Based Medicine (EBM) process to design a novel LLM physician framework, which effectively leverages the retrieval, filtering, and reasoning processes inherent to EBM. Experiments demonstrate that Med-R2 holds superior advantages over existing strategies, while also reducing the substantial computational costs associated with model training. Furthermore, our analysis of model scale and context window size also highlights the scaling capabilities of Med-R2. By scaling the model parameters and adjusting the context window length, we demonstrate the robustness of Med-R2 in achieving optimal performance across a broad spectrum of medical tasks. However, there are some limitations in our work: (1) Due to the professional nature and rapid evolution of medical knowledge, the medical knowledge corpus we construct may not ensure complete coverage of medical information, potentially limiting the upper bound of retrieval precision. (2) During the fine-grained reranking phase, to balance the computational cost and effectiveness, we employ a lightweight proxy model to calculate the usefulness score of the current evidence document, yet it does not fully represent the performance tendencies of the target model during actual evaluating.
@@ Appendix A Evidence-Based Medicine (EBM)
Evidence-Based Medicine (EBM) is defined as the conscientious, explicit, and judicious application of the best current evidence in making decisions regarding the care of individual patients. This practice entails the integration of individual clinical expertise with the most reliable external clinical evidence derived from systematic research [40]. The EBM process typically includes five stages, question formulating, evidence searching, evidence appraising, evidence applying and effect assessing, which aims to make the best possible health care decision through iterative improvements.
To utilize “best evidence”, researchers assess the quality of trials by determining the grading system based on the likelihood that the methods used and the results obtained are less prone to bias and more reliable. The hierarchy of evidence guides the clinical decision-making, since not all evidence is created equal, as described in Figure 6. The evidence hierarchy establishes the priority of references, particularly when conflicting facts are present within the retrieved evidence.
@@
  • Cohort Studies fall into the third-highest category of evidence. Both retrospective and prospective cohort studies are prone to various biases. Prospective cohort studies are considered more reliable, less susceptible to information biases (selection, misclassification, recall), and can establish temporal associations (outcomes following exposure). However, these cohort studies may suffer from confounding biases, which is a major concern that can undermine the validity of their findings.
@@
  • Expert Opinion is considered the lowest level of evidence due to its high susceptibility to bias. Compared to other levels, experts are more likely to choose evidence that confirms their preconceived hypotheses or beliefs, potentially leading to conflicts of interest and a focus on a specific domain while overlooking broader contexts, thereby introducing bias into their perspectives.
Appendix B Method Details
B.1 Question Formulation Details
B.1.1 Details of Query Classification
@@
Evidence-Based Medicine (EBM) Categories  Non-professional queries may fail to clearly articulate the current medical symptoms, thereby hindering the effective retrieval of the necessary evidence. Here we classify medical queries into several types according to the categories of Evidence-Based Medicine (EBM) questions, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc., and then conduct professional medical query reformulations based on their EBM categorization to emphasize specialized retrieval. Instructions for targeted augmentation of clinical queries are outlined in Table 4, where we employ the Qwen2.5-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct model fine-tuned on our 100 manually annotated samples as our question reformulator.
@@
B.1.2 Details of Query Reformulator
@@
Following the instructions outlined in Table 4, we have manually annotated 120 samples (20 samples for each EBM category) to fine-tune Qwen2.5-72B-Instruct, and employed the trained model to professionally reformulate the queries within the medical datasets to enhance the subsequent retrieval of relevant evidence pertaining to the current query from the knowledge corpus.
B.2 Knowledge Corpus Details
B.2.1 Data Sources of Medical Knowledge Base
We organized the medical knowledge corpus from four data sources: academic papers, entries, books, and guidelines. List below are the data sources we built upon:
  • Academic Papers  Academic literature provide valuable insights derived from the latest scientific investigations, which offers a robust theoretical foundation for guiding clinical practice and informing public health decisions. We obtained the publicly available literature from PubMed Central (PMC) [^8] https://pmc.ncbi.nlm.nih.gov/, and exracted valid contents following the processing pipeline of  Hakala et al. [14].
  • Entries  Medical entries provide valuable information across multiple dimensions of healthcare, from direct patient care to cutting-edge research. We extracted and curated medical-related data from the Wikipedia dataset [^9] https://huggingface.co/datasets/wikimedia/wikipedia to build the entries. In the end, we got approximately 470k documents.
  • Books  Medical textbooks are vital resources for medical knowledge retrieval, which can be consulted when faced with complex cases or when seeking updated knowledge about specific conditions. We gathered books from NCBI Bookshelf [^10] https://www.ncbi.nlm.nih.gov/books/, and finally collected 10k documents following the handling process of  Hakala et al. [14].
  • Guidelines  Clinical practice guidelines serve as essential tools in Evidence-Based Medicine (EBM), aiding healthcare providers and patients in making informed decisions regarding diagnosis and treatment. We utilized the guideline data when training MEDITRON series from  Chen et al. [7], and gained approximately 10k documents.
The statistical information on volume and token number across different knowledge resources can be found in Table 1. Based on our analysis, 95.8% of the documents in the entire medical knowledge corpus have a length of less than 10,000 tokens. Consequently, we set the threshold for content segmentation at 10,000 tokens.
B.2.2 Details of the Retrieval Strategy
Motivation for Choosing Hybrid Dense and Sparse Retrieval  The hybrid dense and sparse retrieval leverages the strengths of both dense and sparse representations to enhance retrieval performance, which is particularly beneficial in the medical domain [33]. Dense retrieval (e.g., DPR [26], Sentence-BERT [37]) captures the contextual and deep semantic meanings, while sparse retrieval (e.g., BM25 [38], TF-IDF [36]) excels at accurately matching specific keywords. Such strategy achieves a harmonious integration of semantic comprehension and precise keyword alignment. Prior to the fine-grained reranking process, a substantial number of documents are retrieved and await detailed reordering. Consequently, the goal during the evidence retrieval phase is to identify documents that can assist in addressing the current medical question as comprehensively as possible, which renders hybrid dense and sparse retrieval the most suitable option. In Table 5 we present the strengths and weaknesses of different retrieval strategies:
Table 5: Comparison among different retrieval strategies.
Strategy Semantic Understanding Keyword Precision Recall vs. Precision
Dense Retrieval Strong Weak High recall rate
Sparse Retrieval Weak Strong High precision
Hybrid Dense and Sparse Retrieval Strong Strong Balances both
B.3 Evidence Reranking Details
B.3.1 Components of Fine-Grained Reranker
@@
We detail the scoring clarification and implementation for the fine-grained reranking phase of retrieved documents from Equation 1.
  • Hierarchy of Evidence $f_{h}(x)$: It should be clarified that for the guideline-sourced data retrieved from the knowledge corpus, we assign the $f_{h}$ score directly to 7 ($e=3$) based on the evidence level depicted in Figure 6. It is because the grading system of evidence hierarchy explicitly designates a specific level for guideline data.
  • Usefulness $f_{u}(x)$: 1) Regarding the lightweight proxy models we employed for scoring usefulness $f_{u}$ of the retrieved documents, for the LLaMA series, we used the Sheared-LLaMA-1.3B [48] as a lightweight reference model; for the Qwen series, we utilized Qwen-2.5-1.5B [51]. 2) Moreover, to avoid potential label leakage when computing $f_{u}$ in Equation 3, we have implemented strategies to ensure that the entire framework does not have access to ground truth answers (labels) during the deployment phase, where we employed frontier models’ responses to serve as the “reference answers” for consultation throughout the framework’s operation. Specifically, we utilized the training subsets from MedQA-USMLE, MedQA-MCMLE [21], and MedMCQA [35] to fine-tune the Qwen2.5-72B-Instruct [^11] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct model, resulting in model $M_{F}$. We replaced the labels with responses generated by model $M_{F}$ for the current medical question.
  • General Document Category $f_{g}(x)$: The calculation is based on the categorization and mapping to general document categories, as depicted in Algorithm 1. For this categorization, advanced language models provide probabilities. In addition, the relationship between the general question category and the general document category is multifaceted (illustrated in  Figure 8), which further enhances the robustness of category assignment.
@@
B.3.2 Explanation for Equation Derivation
We have chosen Equation 1 in the form of multiplication instead of additive scoring primarily for the following two reasons:
  • Multi-Variable Synergy: The multiplicative structure emphasizes the synergistic effect, preventing any single variable from dominating. It is necessary for all variables to reach a certain value simultaneously to achieve the expected outcomes. For instance, if $f_{h}(x)$ or $f_{g}(x)$ approaches a lower bound (e.g., $f_{g}(x)\to 0$), even if the usefulness score $f_{u}(x)$ is extremely high, the final outcome $\mathcal{F}(x)$ will still tend towards 0.
  • Metric Self-Adaptation: The value ranges of the variables differ significantly (e.g., $f_{h}(x)\in[1,9]$, $f_{g}(x)\in(0,1)$, $f_{u}(x)\in[0,+\infty)$). Equation 1 built upon the multiplicative formula naturally balances the scales of these variables. In contrast, the weighted summation approach requires manual normalization of variables.
B.4 Evidence Assessment Details
Here, we provide the details of the iterative multi-document retrieval process, as illustrated in Algorithm 2. There are several points that require clarification.
  • Integration of CoT Sequences  It is important to note that during the initial retrieval phase, we did not incorporate the generated CoT sequences as part of the query reformulation $q_{EBM}\left[+q_{CoT}\right]$. Instead, this component was integrated in the second iteration, where we have shown the rationale in Section 5.
  • Accuracy Calculation  To avoid potential label leakage during deployment, we also employed the responses of frontier models to serve as the “reference answers” for consultation throughout the operation (line 9 in Algorithm 2). This implementation is the same as the strategy when computing the usefulness score $f_{u}(x)$ described in Section B.3.
@@ B.4 Evidence Assessment Details
Appendix C Experiments Details
C.1 Details of Medical Datasets
@@
  • MedQA-USMLE and MedQA-MCMLE [21]: The MedQA dataset is derived from the professional medical board exams, covering languages of English, simplified Chinese, and traditional Chinese. In this study, we employ the subsets collected from US and Mainland China. The MedQA-USMLE dataset includes text materials from 18 English medical textbooks, while the MedQA-MCMLE dataset is constructed from 33 simplified Chinese medical textbooks.
  • MedMCQA [35]: The MedMCQA dataset is a large-scale multi-choice question answering data corpus that covers 2,400 healthcare topics and 21 medical subjects. Topics of MedMCQA span from medicine (endocrinology, infection, haematology, respiratory, etc.), surgery (general surgery, endocrinology, breast, and vascular surgery, etc.) to radiology & biochemistry, which are sourced from both real-world scenarios and simulated examinations.
  • PubMedQA [22]: The PubMedQA dataset is a novel biomedical question answering dataset originated from PubMed abstracts. To ensure equity, we exclude the origin-given contexts during the retrieval phase. Here we choose the 1,000 expert-annotated instances for downstream task evaluation. PubMedQA is included only in our test set due to the limited number of instances.
  • MMLU-Med [16]: MMLU-Med is a set of six medical tasks (anatomy, clinical knowledge, professional medicine, human genetics, college medicine, college biology) derived from the MMLU dataset. MMLU-Med is also included only in our test set due to its limited quanity.
@@
MedQA-USMLE [21] 10178 2535 10 12723
MedQA-MCMLE [21] 27400 6840 11 34251
MedMCQA [35] 182822 4170 13 187005
PubMedQA [22] - 990 10 1000
MMLU-Med [16] - 1080 9 1089
C.2 Details of Baselines
  • Frontier Models: Frontier models are considered to represent the pinnacle of performance across various dimensions of LLMs, serving as the strongest baselines. Here, we have selected several state-of-the-art LLMs to establish the maximum potential of model performance on several medical benchmarks, including GPT-4o [17], Claude3.5-Sonnet [3], and DeepSeek-V3 [30].
  • Open-Sourced Medical Models: These models stand for the domain-specific models that were specifically trained on medical data. We have selected PMC-LLaMA-7B, PMC-LLaMA-13B [47], MEDITRON-7B and MEDITRON-70B [7] to represent the open-sourced medical models for comparison to assess Med-R2’s relative advantage compared to specialized medical AI systems.
  • Direct Response: In this case, we employ the base model to directly respond to medical queries without aid of additional training on any dataset or augmentation from external knowledge bases.
  • Vanilla RAG [28]: It represents the most traditional and fundamental strategy for utilizing external knowledge bases to assist models in answering questions. In this study, we employ the same medical knowledge base as Med-R2 , but directly utilizing the raw text retrieved by FAISS and combining it with the original query to test the model’s performance on medical tasks.